OpenStudy (anonymous):

Find the slope of the line when the points are (9,2) and (4,0)

OpenStudy (saifoo.khan):

\[\frac{0-2}{4-9}\]

OpenStudy (saifoo.khan):

\[\frac{-2}{-5} \]

OpenStudy (saifoo.khan):

therefore: \[m = \frac{2}{5}\]